# Architecting mechanisms of damage in topological metamaterials

### v1

@@ Introduction
In this section we focus on how the form of FM introduced at the pre-crack edge influences the behaviour and damage evolution for mode-I opening. Initially, we focus on the results for RK, LTK, LP1, RP1 and LUP1, for which $(x_{\mathrm{1}},x_{\mathrm{2}},x_{\mathrm{3}})$ are (0,0,0), (0.1,0.1,0.1), (-0.1,0.1,0.1), (0.1,-0.1,-0.1) and (-0.1,0.1,-0.1) respectively. The results for the remaining equivalence class, RDP, are provided in the SI Appendix, along with RTK. Additionally, we include RP and LP lattices, which have chiral unit cells, i.e. mirror-symmetric with respect to the crack plane. Results for achiral unit-cells from the same classes are also found in the SI Appendix.
Dominant Lattice Behavior
@@
In practice, bonds are constrained from freely rotating around each site due to the local bending resistance associated with finite thickness joints. However, contrasting the idealized GH modes (i.e., affine strain-producing mechanisms) with the lattice deformation in the simulator, offers insights into the mechanics while traversing SR-space. Geometries that would ideally possess a GH mode will, in reality, become bending-dominated at high SR values, if the applied boundary displacements closely match those required by the GH mode (e.g., Fig. 1c). The LTK, LP, and RP lattices, all belong in this category. On the contrary, lattices that either do not posses a GH mode, such as the RK lattice, or whose GH mode does not produce boundary displacements compatible with the simulated ones, such as LUP (Fig. 1d), remain stretching-dominated throughout the SR range. Here, there is a caveat for the RK lattice, whose uniform FM depicted in Fig. 1b is not classified as a GH mode, as it does not produce any affine strain to linear order in the displacements. This mode is rather an infinitesimal increment of the GH mode [30], which can only map to boundary displacements in compression through an instability, leading to stretching-dominated behaviour in tension, across the SR range.
@@
Figure 1: (a) Axial energy portion in the entire lattice preceding the first failure event. Visualisation of the affine strain producing mechanism (GH mode) for: (b) RK; (c) LP1; and, (d) LUP1. Comparison of the internal angle of a unit cell from the simulations ($\alpha_{\mathrm{s}}$) with that required by the GH mode ($\alpha_{\mathrm{GH}}$) to displace the boundary the equivalent distance (i.e. $\alpha_{\mathrm{d}}=\alpha_{\mathrm{s}}-\alpha_{\mathrm{GH}}$) for: (e,h) LTK, (f,i) LP1; and, (g,j) RP1 at SR = 3 and 30 respectively. The difference in angle is visualised through the colour directly to the right of each unit cell.
@@
To explore whether non-idealized lattices deform along the GH mode, further visual comparisons are presented for low (Fig. 1e-g) and high (Fig. 1h-j) SR. These figures highlight the difference between the simulated local angle, $\alpha_{\mathrm{s}}$, and the expected GH angle, $\alpha_{\mathrm{GH}}$. Despite the majority of the bulk following the GH mode for the bending-dominated lattices at $\text{SR}=30$, they also posses regions that diverge from this behaviour. Bending stresses mirror this trend (Fig. 2a,c,l), with higher stress values appearing in the same regions as the maps for $\alpha_{\mathrm{d}}$ in Fig. 1. Below, we argue that the form of the regions that diverge from the GH mode can be connected to the edge actuation through the bulk-boundary correspondence. Thus, the phenomena observed can be interpreted in light of the compliance added to the system by the pre-crack.
The bending stresses along the mid-plane (Fig. 2i) reveal the FM actuation both at the crack edge and within the lattice. For LTK and LP1, the actuation is mainly concentrated in a sharp peak at the crack tip, while for RP1, the peak remains close to the uniform value observed in the bulk. All lattices are relaxed along the rest of the crack edge, to the left of the crack tip. By modeling the FM as a modulation of the uniform zero-energy GH mode [33], without restricting the description to linear order, we can leverage bulk-boundary correspondence, prescribing an analogue of the observed dominant actuation pattern along the mid-plane, to visualize the resulting FM amplitude, $|\phi|$, in the bulk. In particular, we prescribe a Gaussian spike of unit height and width, centred at the crack-tip, for the LP/LTK lattices (Fig. 2e-h), and a sigmoid function centred at the crack-tip for RP (Fig. 2m,o). Different prescriptions can be superimposed, so one could easily add the Gaussian and sigmoid solutions together with different relative amplitudes, to reach a more accurate visualisations of the FM in the bulk for each lattice. However, the goal here is to capture the most dominant components of the disturbance; the method should be seen as a simple heuristic (full details of the visualisation can be found in SI Appendix).
@@
Figure 2: Maximum normalised bending stress in the elements cross section $\sigma_{\mathrm{\textbf{Bend}}}=|\sigma_{\mathrm{B}}/\sigma_{\mathrm{B}}^{\mathrm{u}}|$ and visualisation of FM through the magnitude of the total modulation ($|\phi|$) of the wave spectrum for: (a,e) RK (0.1,0.1,0.1); (b,f) LP2 (-0.02,0.18,0.18); (c,g) LP1 (-0.1,0.1,0.1); (d,h) LP3 (-0.16,0.04,0.04); (l,m) RP1 (0.1,-0.1,-0.1); and, (n,o) RP2 (0.02,-0.18,-0.18). (i) Normalised bending stress along the arc length. The bending stress is normalised with the peak bending stress along this length. The arc length is illustrated in (a). (j) Decay spectrum of zero modes on a horizontal surface; and, (k) associated wave angles. All figures associated with simulations are for $\mathrm{SR}=30$.
@@
Three additional lattices within the LP/RP ($\mp$,$\pm$,$\pm$) equivalence classes are simulated to further illustrate how the form of the FM influence the stress profile within the lattice. Even with the simplified aforementioned prescriptions, the FM fields in the bulk (Fig. 2e-h,m,o) map qualitatively well onto the major components of the bending stresses in the lattice (Fig. 2a-d,l,n). To better understand the form of the FM in the bulk, Fig. 2j-k, can offer valuable insights. Firstly, the penetration into the bulk of the localised disturbance can be connected to the decay rate $\kappa^{\mathrm{FM}}(q_{\mathrm{1}})$ spectrum. At the homogenisation limit (i.e small $q_{\mathrm{1}}$), non-polarised lattices (e.g LTK) are expected to decay faster because they have $\kappa^{\mathrm{FM}}(q_{\mathrm{1}})\sim\mathcal{O}(q_{\mathrm{1}})$, compared to $\kappa^{\mathrm{FM}}(q_{\mathrm{1}})\sim\mathcal{O}(q^{2}_{\mathrm{1}})$ for polarized lattices [20]. Here, we observe that, in general, the magnitude of $\kappa^{\mathrm{FM}}(q_{\mathrm{1}})$ across the whole range is more indicative, which makes sense for a localised disturbance, as it samples a wide spectrum in $q_{\mathrm{1}}$-space. Subsequently, LTK and LP2 decay quickly into the bulk, with the deformation away from the crack-tip reverting back to that associated with the idealised GH mode. Contrarily, the spectra of LP1 and LP3 exhibit lower decay rates, with the disturbances influencing the behaviour deeper into the bulk.
@@
Dominant Lattice Failure
@@
Figure 3: (a) Axial stress fraction in member that fractures preceding the first failure event. (b) The damage process is visualised with a line traversing the domain, connecting separated nodes (fracture locations) within the lattice, indicated by yellow markers. For clarity we do not show the microstructure throughout the domain. The damage process for SR = 3, 6, 10, 30 are plotted with green, purple, black, and blue lines and associated rectangular markers respectively (refer to (a)). Arrows indicate the direction of damage and circular markers indicate a discontinuity in the damage path. Damage paths for: (c) RK; (d) LTK; (e) LP1; (f) RP1; and, (g) LUP1.
While the LUP1 global behaviour is stretching-dominated, the fracture of its elements is strongly influenced by the bending (Fig. 3a) associated with the FM localising at the pre-crack edge at large SR (Fig. 4a,d). As such, strong alignment of the damage path with the long-wavelength limit of $\theta_{\mathrm{2}}$ (Fig. 4i) is observed. Simulating other lattices from the equivalence class (-,+,-) at SR = 30, it indicates this is a robust phenomenon: for LUP2 (-0.18,0.04,-0.1) and LUP3 (-0.16,0.04,-0.16) the damage path (Fig. 4g) follows the long-wavelength limit of $\theta_{2}$ and $\theta_{1}$ respectively (Fig. 4d-f,i). This results in the damage evolution following a shallower angle than LUP1. Close to the long-wavelength limit, $\theta_{1}$ propagates further into the bulk for LUP1-3. While the lattice appears to favour propagating the crack in the direction with the smallest $\kappa^{\mathrm{FM}}$ (i.e. LUP3), if $\theta$ is behind the crack tip, it instead favours the modes that propagate forward, ahead of the crack tip (i.e. LUP1 and 2). This phenomenon is also observed for the RDP lattices, provided in the SI Appendix. For lattices in the ($\pm$$\mp$,$\pm$) equivalence classes, at large SR, there is a strong alignment of the damage path with the FM localised at the crack tip, providing a control parameter to steer the localised fracture of its underlying microstructure. As the SR reduces, the crack tends towards a horizontal path (Fig. 3g), with these FM having a reduced influence on the damage.
Figure 4: Maximum normalised bending stress in the elements cross section $\sigma_{\mathrm{\textbf{Bend}}}=|\sigma_{\mathrm{B}}/\sigma_{\mathrm{B}}^{\mathrm{u}}|$ and visualisation of FM through the magnitude of the total modulation ($|\phi|$) of the wave spectrum for: (a,d) LUP1 (-0.1,0.1,-0.1); (b,e) LUP2 (-0.18,0.04,-0.1); and, (c,f) LUP3 (-0.16,0.04,-0.16). (g) Damage path for LUP1, LUP2, LUP3 at SR = 30. (h) Damage path for lattice with a domain wall separating RDP1 and LUP1. (i) $\theta$ spectrum for LUP1, LUP2 and LUP3. All figures associated with simulations are for SR = 30.
Effect of Discontinuities
@@
Topological domain walls
@@
Figure 5: Normalised applied stress vs nominal strain for: (a) SR = 3; and, (b) SR = 30. Maximum normalised stress in the elements cross section for D-FM, D-SSS1 and D-SSS2 at specified nominal strains for: (i,iii) SR = 3; and, (ii,iv-viii) SR = 30. Where $\sigma_{\mathrm{M}}=\sigma_{\mathrm{A}}/\sigma_{\mathrm{A}}^{\mathrm{u}}\pm|\sigma_{\mathrm{B}}/\sigma_{\mathrm{B}}^{\mathrm{u}}|$. In (v), (vi), and (viii) yellow stars indicated nodes that are added into the system due to the separation of elements (element fracture) along the domain wall. For clarity, an insert is also included where circular yellow markers indicate these nodes.
@@
Figure 6: Maximum normalised stress in the elements cross section for (a) LP1 and (b) LUP1 at SR = 30. Damage path with pinned boundary for: (c) LP1 at SR = 10 and 30; (d) LP1 at SR = 30 with 10 unit pre-crack length; (e) LP2 at SR =30; (f) LUP1 at SR = 30; and, (g) LUP1 at SR = 30 with 50 units along the length. In (c-g) we also highlight the zones where the SSS at the long-wavelength limit may develop with arrows either side of these zones. If the SSS can’t develop, we only plot the direction of the SSS with a single arrow. Maximum normalised bending stress in the elements cross section at SR = 30 for: (h) LP1 with 10 unit pre-crack length; (i) LUP1; and, (j) LUP1 with 50 units along the length.
@@
The lattice size should not significantly affect its dominant behaviour when the SSS at the long-wavelength limit dominate the mechanical response. This occurs when $l_{\mathrm{\theta}}$ is above a critical length that localises the SSS towards this limit, as these modes do not decay. In SI Appendix we simulate a series of LP1 lattices with increasing domain size but identical aspect ratio. Keeping the aspect ratio constant ensures the SSS zones relative size remains constant. Plotting the axial energy portion for these lattices shows that the dominant behaviour does not vary significantly, and therefore, for this aspect ratio the SSS associated with the long-wavelength limit dominates the response.
@@
A similar damage evolution is observed in other confined non-topologically polarised lattices when $z>2d$ [34]. In these lattices, an incomplete debonding type failure occurs, leaving bridging elements along the interface. Subsequently, these elements fracture, completely debonding the interface. SSS are expected to form within the lattice as it is confined and $z>2d$. Therefore, this initial failure process may occur due to the removal of these SSS. Once removed, the remaining bridging elements become bending-dominated (i.e. the local $z\leq 2d$), leading to the observed toughening behaviour. For lattices with $z<2d$ this bridging is not as readily observed [34].
When the size of the lattice prevents the SSS from developing, the dominant behaviour and failure do not vary greatly between roller or pinned boundary conditions. This is the case for LP2, where the damage path is identical for both boundary conditions (Fig. 6e) and the axial energy only rises from $3\%$ to $11\%$. LP2 would require a minimum length of 49 units for the SSS at the long-wavelength limit to develop. Increasing the lattice length to 50 units results in these SSS developing, and in SI Appendix we show that the fracture initiates at the top and bottom edges. The axial energy portion further increases to $22\%$, highlighting that the original length of LP2 was not large enough for these SSS to develop.
For LUP1 the axial energy portion goes from approximately 70% to 95% when the boundary conditions are switched to pinned for SR = 30. The SSS long-wavelength limit zones shown in Fig. 6f strongly correlate to the stress that develops within the lattice (Fig. 6b). The SSS corresponding to $\theta_{1}$ and $\theta_{2}$ require 2 and 34 units along the length to develop for the domain size simulated. The SSS associated with $\theta_{2}$, therefore, does not occur. The change in the stress profile with respect to the form of these modes is provided in SI Appendix for LUP1-3.
@@
Conclusions
Designing the mechanisms of damage in metamaterial has predominately leaned on the inclusion of heterogeneities to overcome the plethora of influencing parameters and competing length scales. In this work, we instead show that homogeneous Maxwell lattices have topological and geometry-dependent parameters that dominate the locations of stress concentrations, particularly at large SR. These parameters are derived from the form of the FM and SSS that exponentially localise on bulk discontinuities. The localisation of these modes is captured through a topological index, defining the equivalence class and constraining the microstructure to those which exhibit an identical topological protected directional preference. Within each equivalence class, the form of these modes i.e. the decay and angle of propagation, can be tuned by distorting the microstructure without affecting its topological polarisation $\textbf{P}_{\mathrm{T}}$. Using a beam element simulator, we numerically probed the effect of these modes over a range of slenderness for a set of distorted kagome lattices that span the possible equivalence classes.

### v2

@@ 4 Behaviour of Topologically Distinct Lattices
In this section we focus on how the form of FM introduced at the pre-crack edge influences the behaviour and damage evolution for mode-I opening. Initially, we focus on the results for RK, LTK, LP1, RP1 and LUP1, for which $(x_{\mathrm{1}},x_{\mathrm{2}},x_{\mathrm{3}})$ are (0,0,0), (0.1,0.1,0.1), (-0.1,0.1,0.1), (0.1,-0.1,-0.1) and (-0.1,0.1,-0.1) respectively. The results for the remaining equivalence class, RDP, are provided in the C, along with RTK. Additionally, we include RP and LP lattices, which have chiral unit cells, i.e. mirror-symmetric with respect to the crack plane. Results for achiral unit-cells from the same classes are also found in the E.
4.1 Dominant Lattice behaviour
@@
Figure 4: Dominant behaviour of lattices. (a) Axial energy portion in the entire lattice preceding the first failure event. Visualisation of the affine strain producing mechanism (GH mode) for: (b) RK; (c) LP1; and, (d) LUP1. Comparison of the internal angle of a unit cell from the simulations ($\alpha_{\mathrm{s}}$) with that required by the GH mode ($\alpha_{\mathrm{GH}}$) to displace the boundary the equivalent distance (i.e. $\alpha_{\mathrm{d}}=\alpha_{\mathrm{s}}-\alpha_{\mathrm{GH}}$) for: (e, h) LTK, (f, i) LP1; and, (g, j) RP1 at SR = 3 and 30 respectively. The difference in angle is visualised through the colour directly to the right of each unit cell.
In practice, bonds are constrained from freely rotating around each site due to the local bending resistance associated with finite thickness joints. However, contrasting the idealised GH modes (i.e., affine strain-producing mechanisms) with the lattice deformation in the simulator, offers insights into the mechanics while traversing SR-space. Geometries that would ideally possess a GH mode will, in reality, become bending-dominated at high SR values, if the applied boundary displacements closely match those required by the GH mode (e.g., Fig. 4c). The LTK, LP, and RP lattices, all belong in this category. On the contrary, lattices that either do not posses a GH mode, such as the RK lattice, or whose GH mode does not produce boundary displacements compatible with the simulated ones, such as LUP (Fig. 4d), remain stretching-dominated throughout the SR range. Here, there is a caveat for the RK lattice, whose uniform FM depicted in Fig. 4b is not classified as a GH mode, as it does not produce any affine strain to linear order in the displacements. This mode is rather an infinitesimal increment of the GH mode (Hutchinson and Fleck, 2006), which can only map to boundary displacements in compression through an instability, leading to stretching-dominated behaviour in tension, across the SR range.
@@
To explore whether non-idealised lattices deform along the GH mode, further visual comparisons are presented for low (Fig. 4e-g) and high (Fig. 4h-j) SR. These figures highlight the difference between the simulated local angle, $\alpha_{\mathrm{s}}$, and the expected GH angle, $\alpha_{\mathrm{GH}}$. Despite the majority of the bulk following the GH mode for the bending-dominated lattices at $\text{SR}=30$, they also posses regions that diverge from this behaviour. Bending stresses mirror this trend (Fig. 5a,c,l), with higher stress values appearing in the same regions as the maps for $\alpha_{\mathrm{d}}$ in Fig. 4. Below, we argue that the form of the regions that diverge from the GH mode can be connected to the edge actuation through the bulk-boundary correspondence. Thus, the phenomena observed can be interpreted in light of the compliance added to the system by the pre-crack.
@@
Figure 5: Mapping of bending stresses through the bulk-boundary correspondence. Maximum normalised bending stress in the elements cross section $\sigma_{\mathrm{\textbf{Bend}}}=|\sigma_{\mathrm{B}}/\sigma_{\mathrm{B}}^{\mathrm{u}}|$ and visualisation of FM through the magnitude of the total modulation ($|\phi|$) of the wave spectrum for: (a, e) RK (0.1,0.1,0.1); (b, f) LP2 (-0.02,0.18,0.18); (c, g) LP1 (-0.1,0.1,0.1); (d, h) LP3 (-0.16,0.04,0.04); (l, m) RP1 (0.1,-0.1,-0.1); and, (n, o) RP2 (0.02,-0.18,-0.18). (i) Normalised bending stress along the arc length. The bending stress is normalised with the peak bending stress along this length. The arc length is illustrated in (a). (j) Decay spectrum of zero modes on a horizontal surface; and, (k) associated wave angles. All figures associated with simulations are for SR = 30.
The bending stresses along the mid-plane (Fig. 5i) reveal the FM actuation both at the crack edge and within the lattice. For LTK and LP1, the actuation is mainly concentrated in a sharp peak at the crack tip, while for RP1, the peak remains close to the uniform value observed in the bulk. All lattices are relaxed along the rest of the crack edge, to the left of the crack tip. By modeling the FM as a modulation of the uniform zero-energy GH mode (Rocklin et al., 2017), without restricting the description to linear order, we can leverage bulk-boundary correspondence, prescribing an analogue of the observed dominant actuation pattern along the mid-plane, to visualize the resulting FM amplitude, $|\phi|$, in the bulk. In particular, we prescribe a Gaussian spike of unit height and width, centred at the crack-tip, for the LP/LTK lattices (Fig. 5e-h), and a sigmoid function centred at the crack-tip for RP (Fig. 5m,o). Different prescriptions can be superimposed, so one could easily add the Gaussian and sigmoid solutions together with different relative amplitudes, to reach a more accurate visualisations of the FM in the bulk for each lattice. However, the goal here is to capture the most dominant components of the disturbance; the method should be seen as a simple heuristic (further details of the visualisation can be found in B).
@@
Three additional lattices within the LP/RP ($\mp$,$\pm$,$\pm$) equivalence classes are simulated to further illustrate how the form of the FM influence the stress profile within the lattice. Even with the simplified aforementioned prescriptions, the FM fields in the bulk (Fig. 5e-h,m,o) map qualitatively well onto the major components of the bending stresses in the lattice (Fig. 5a-d,l,n). To better understand the form of the FM in the bulk, Fig. 5j-k, can offer valuable insights. Firstly, the penetration into the bulk of the localised disturbance can be connected to the decay rate $\kappa^{\mathrm{FM}}(q_{\mathrm{1}})$ spectrum. At the homogenisation limit (i.e small $q_{\mathrm{1}}$), non-polarised lattices (e.g LTK) are expected to decay faster because they have $\kappa^{\mathrm{FM}}(q_{\mathrm{1}})\sim\mathcal{O}(q_{\mathrm{1}})$, compared to $\kappa^{\mathrm{FM}}(q_{\mathrm{1}})\sim\mathcal{O}(q^{2}_{\mathrm{1}})$ for polarised lattices (Kane and Lubensky, 2014). Here, we observe that, in general, the magnitude of $\kappa^{\mathrm{FM}}(q_{\mathrm{1}})$ across the whole range is more indicative, which makes sense for a localised disturbance, as it samples a wide spectrum in $q_{\mathrm{1}}$-space. Subsequently, LTK and LP2 decay quickly into the bulk, with the deformation away from the crack-tip reverting back to that associated with the idealised GH mode. Contrarily, the spectra of LP1 and LP3 exhibit lower decay rates, with the disturbances influencing the behaviour deeper into the bulk.
@@
4.2 Dominant Lattice Failure
@@
Figure 6: Dominant failure of lattices. (a) Axial stress fraction in member that fractures preceding the first failure event. (b) The damage process is visualised with a line traversing the domain, connecting separated nodes (fracture locations) within the lattice, indicated by yellow markers. For clarity we do not show the microstructure throughout the domain. The damage process for SR = 3, 6, 10, 30 are plotted with green, purple, black, and blue lines and associated rectangular markers respectively (refer to (a)). Arrows indicate the direction of damage and circular markers indicate a discontinuity in the damage path. Damage paths for: (c) RK; (d) LTK; (e) LP1; (f) RP1; and, (g) LUP1.
While the LUP1 global behaviour is stretching-dominated, the fracture of its elements is strongly influenced by the bending (Fig. 6a) associated with the FM localising at the pre-crack edge at large SR (Fig. 7a,d). As such, strong alignment of the damage path with the long-wavelength limit of $\theta_{\mathrm{2}}$ (Fig. 7i) is observed. Simulating other lattices from the equivalence class (-,+,-) at SR = 30, it indicates this is a robust phenomenon: for LUP2 (-0.18,0.04,-0.1) and LUP3 (-0.16,0.04,-0.16) the damage path (Fig. 7g) follows the long-wavelength limit of $\theta_{2}$ and $\theta_{1}$ respectively (Fig. 7d-f,i). This results in the damage evolution following a shallower angle than LUP1. Close to the long-wavelength limit, $\theta_{1}$ propagates further into the bulk for LUP1-3. While the lattice appears to favour propagating the crack in the direction with the smallest $\kappa^{\mathrm{FM}}$ (i.e. LUP3), if $\theta$ is behind the crack tip, it instead favours the modes that propagate forward, ahead of the crack tip (i.e. LUP1 and 2). This phenomenon is also observed for the RDP lattices, provided in the C. For lattices in the ($\pm$$\mp$,$\pm$) equivalence classes, at large SR, there is a strong alignment of the damage path with the FM localised at the crack tip, providing a control parameter to steer the localised fracture of its underlying microstructure. As the SR reduces, the crack tends towards a horizontal path (Fig. 6g), with these FM having a reduced influence on the damage.
Figure 7: Damage path control using FM. Maximum normalised bending stress in the elements cross section $\sigma_{\mathrm{\textbf{Bend}}}=|\sigma_{\mathrm{B}}/\sigma_{\mathrm{B}}^{\mathrm{u}}|$ and visualisation of FM through the magnitude of the total modulation ($|\phi|$) of the wave spectrum for: (a, d) LUP1 (-0.1,0.1,-0.1); (b, e) LUP2 (-0.18,0.04,-0.1); and, (c, f) LUP3 (-0.16,0.04,-0.16). (g) Damage path for LUP1, LUP2, LUP3 at SR = 30. (h) Damage path for lattice with a domain wall separating RDP1 and LUP1. (i) $\theta$ spectrum for LUP1, LUP2 and LUP3. All figures associated with simulations are for SR = 30.
5 Effect of Discontinuities
@@
5.1 Topological domain walls
@@
Figure 8: Effect of domain walls. Normalised applied stress vs nominal strain for: (a) SR = 3; and, (b) SR = 30. Maximum normalised stress in the elements cross section for D-FM, D-SSS1 and D-SSS2 at specified nominal strains for: (i,iii) SR = 3; and, (ii,iv-viii) SR = 30. Where $\sigma_{\mathrm{M}}=\sigma_{\mathrm{A}}/\sigma_{\mathrm{A}}^{\mathrm{u}}\pm|\sigma_{\mathrm{B}}/\sigma_{\mathrm{B}}^{\mathrm{u}}|$. In (v), (vi), and (viii) yellow stars indicated nodes that are added into the system due to the separation of elements (element fracture) along the domain wall. For clarity, an insert is also included where circular yellow markers indicate these nodes.
@@ 5.2 Additional constraints
Figure 9: Effect of pinned boundary conditions. Maximum normalised stress in the elements cross section for (a) LP1 and (b) LUP1 at SR = 30. Damage path with pinned boundary for: (c) LP1 at SR = 10 and 30; (d) LP1 at SR = 30 with 10 unit pre-crack length; (e) LP2 at SR =30; (f) LUP1 at SR = 30; and, (g) LUP1 at SR = 30 with 50 units along the length. In (c-g) we also highlight the zones where the SSS at the long-wavelength limit may develop with arrows either side of these zones. If the SSS can’t develop, we only plot the direction of the SSS with a single arrow. Maximum normalised bending stress in the elements cross section at SR = 30 for: (h) LP1 with 10 unit pre-crack length; (i) LUP1; and, (j) LUP1 with 50 units along the length.
@@
The lattice size should not significantly affect its dominant behaviour when the SSS at the long-wavelength limit dominate the mechanical response. This occurs when $l_{\mathrm{\theta}}$ is above a critical length that localises the SSS towards this limit, as these modes do not decay. In D.1 we simulate a series of LP1 lattices with increasing domain size but identical aspect ratio. Keeping the aspect ratio constant ensures the SSS zones relative size remains constant. Plotting the axial energy portion for these lattices shows that the dominant behaviour does not vary significantly, and therefore, for this aspect ratio the SSS associated with the long-wavelength limit dominates the response.
@@
A similar damage evolution is observed in other confined non-topologically polarised lattices when $z>2d$ (Hedvard et al., 2024). In these lattices, an incomplete debonding type failure occurs, leaving bridging elements along the interface. Subsequently, these elements fracture, completely debonding the interface. SSS are expected to form within the lattice as it is confined and $z>2d$. Therefore, this initial failure process may occur due to the removal of these SSS. Once removed, the remaining bridging elements become bending-dominated (i.e. the local $z\leq 2d$), leading to the observed toughening behaviour. For lattices with $z<2d$ this bridging is not as readily observed (Hedvard et al., 2024).
When the size of the lattice prevents the SSS from developing, the dominant behaviour and failure do not vary greatly between roller or pinned boundary conditions. This is the case for LP2, where the damage path is identical for both boundary conditions (Fig. 9e) and the axial energy only rises from $3\%$ to $11\%$. LP2 would require a minimum length of 49 units for the SSS at the long-wavelength limit to develop. Increasing the lattice length to 50 units results in these SSS developing, and in D.2 we show that the fracture initiates at the top and bottom edges. The axial energy portion further increases to $22\%$, highlighting that the original length of LP2 was not large enough for these SSS to develop.
For LUP1 the axial energy portion goes from approximately 70% to 95% when the boundary conditions are switched to pinned for SR = 30. The SSS long-wavelength limit zones shown in Fig. 9f strongly correlate to the stress that develops within the lattice (Fig. 9b). The SSS corresponding to $\theta_{1}$ and $\theta_{2}$ require 2 and 34 units along the length to develop for the domain size simulated. The SSS associated with $\theta_{2}$, therefore, does not occur. The change in the stress profile with respect to the form of these modes is provided in the D.3 for LUP1-3.
@@
6 Conclusions
Designing the mechanisms of damage in metamaterials has predominately leaned on the inclusion of heterogeneities to overcome the plethora of influencing parameters and competing length scales. In this work, we instead show that homogeneous Maxwell lattices have topological and geometry-dependent parameters that dominate the locations of stress concentrations, particularly at large SR. These parameters are derived from the form of the FM and SSS that exponentially localise on bulk discontinuities. The localisation of these modes is captured through a topological index, defining the equivalence class and constraining the microstructure to those which exhibit an identical topological protected directional preference. Within each equivalence class, the form of these modes i.e. the decay and angle of propagation, can be tuned by distorting the microstructure without affecting its topological polarisation $\textbf{P}_{\mathrm{T}}$. Using a beam element simulator, we numerically probed the effect of these modes over a range of slenderness for a set of distorted kagome lattices that span the possible equivalence classes.